OpenStudy (anonymous):

find an equation of the line containing the given pair of points (-2,-9) (-6,-7)

OpenStudy (anonymous):

y=(1/2)x-4

OpenStudy (anonymous):

okay, i am not good at math, could you please tell me how you got that answer :D

OpenStudy (anonymous):

A line is linear. y=mx+b If a line passes through a certain point, that line's equation has to be true at that point; it has to be true for those specific x and y values. You can simply substitute the x and y values into the equation of your line. (-2,-9) means x = -2 and y= -9 y=mx+b -9= -2m+b Now substitute in the x and y values from your other point into the equation y=mx+b, to get a second equation with actual numbers in it. y=mx+b x= -6 y = -7 -7= -6m+b so now you have two equations with actual numbers in them... -9= -2m+b and -7= -6m+b now, you want to figure out what m and b are. There are multiple ways to do this. One way is to solve one equation for either m or b, like this: -9= -2m+b -9+2m= b Then, substitute what b equals (shown above) into the other equation: -7= -6m+b b = (-9+2m) -7= -6m+(-9+2m) -7= -4m-9 2 = -4m 2/(-4) = m m = -1/2 once you've found the value r m it's pretty simple to find out what b is, or vis versa. Just take either of your two equations, -9= -2m+b OR -7= -6m+b and replace m with "-1/2" -9= -2(-1/2)+b OR -7= -6(-1/2)+b and solve either one of them for b. y= mx+b once you've found m and b, you can write your line's equation: y= (-1/2)x-10